solve for x
(9/4x)+(1/3)=(5/2x)
I don't even know where to begin

Answer by ilana(307)

Are you sure you don't mean (9/4)x+(1/3)=(5/2)x? The way you wrote it says x is in the denominator, with 4 and 2. I am guessing that's not the case. If it is, repost and I'm terribly sorry. I am going to answer my version because I like it better.
I would begin by getting rid of all of those evil fractions! There is a 4, a 3, and a 2, so I need to multiply by something to cancel out all 3 of those. 12 is the smallest number that will do that.
12
Distribute the 12.
12(9/4)x+12(1/3)=12(5/2)x
3(9)x+4(1)=6(5)x
27x+4=30x Yay, no more fractions! (for now)
4=3x
x=4/3